OpenStudy (anonymous):

please help if 4 over 5 ton of concrete covers 5 over 6 of a bridge, how many tons of concrete are required to cover the entire bridge? 1 over 5 ton 9 over 30 ton 2 over 3 ton 24 over 25 ton

OpenStudy (anonymous):

4/5 : 5 /6 = x :1 just solve proportion you'll get 5x/6 =4/5

OpenStudy (anonymous):

solve for x and you will get answer, do you get it? :)

OpenStudy (anonymous):

.8

OpenStudy (anonymous):

lets try this together, you have 5x/6 =4/5 you get 5/6 on other side 5x/6 =4/5 // *6 5x = 4*6/5 // :5 x = (4*6)/(5*5)

OpenStudy (anonymous):

do you know the answer now?

OpenStudy (anonymous):

.96

OpenStudy (anonymous):

D

OpenStudy (anonymous):

you don't need to divide them :) yes, it's D well done

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're wellcome :)

OpenStudy (anonymous):

how do i give meadals to people

OpenStudy (anonymous):

hahah i have no idea, i'm new here as well :D

OpenStudy (anonymous):

i gave you one i think

OpenStudy (anonymous):

i gave you one i think